OpenStudy (anonymous):

help !!! Given the facts that f(9) = 11 and f'(9) = 3, use the difference quotient to estimate the value of f( 9.5 ). f(9.5) ~=

OpenStudy (anonymous):

\[f^2=2f*f'\]

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

do I plug 9.5 for f?

OpenStudy (anonymous):

do I plug 9.5 for f?

OpenStudy (turingtest):

\[f'(x)\approx\frac{f(x+\Delta x)-f(x)}{\Delta x}\]\[f(x+\Delta x)\approx f(x)+f'(x)\Delta x\]\[f(9.5)=f(9+0.5)\approx f(9)+f'(9)(0.5)\]

OpenStudy (turingtest):

(you plug in 0.5 for delta x is the deal)

OpenStudy (anonymous):

ok so I just plug in 0.5 and multiply with 9 and I get the answer

OpenStudy (anonymous):

right?

OpenStudy (turingtest):

? what do you think the answer is?

OpenStudy (anonymous):

4.5

OpenStudy (turingtest):

no what is f'(9) ?

OpenStudy (anonymous):

0

OpenStudy (turingtest):

why do you think it's zero no, it's given in the question

OpenStudy (anonymous):

answer is 9.5?

OpenStudy (turingtest):

no, how are you getting these answers?

OpenStudy (anonymous):

I thought I was trying to get the derivitave of f'(9)

OpenStudy (turingtest):

f'(9) is a constant, so its derivative would be zero we don't need that, we just need the value of f(9), f'(9) and Delta x the first two are given in the problem, and I already told you the value of Delta x

OpenStudy (anonymous):

so 0.5 x

OpenStudy (anonymous):

and then solve

OpenStudy (turingtest):

I don't know what you mean, show your work please

OpenStudy (anonymous):

what I mean is that would put 0.5 for delta then solve w/ x

OpenStudy (turingtest):

If you knew how to solve the problem you would have done it rather than explain it, it takes just a moment since I already gave you the formula above Again, starting with the difference quotient:\[f'(x)\approx\frac{f(x+\Delta x)-f(x)}{\Delta x}\]rearranging:\[f(x+\Delta x)\approx f(x)+f'(x)\Delta x\]sub in for a particular point, and not that the difference between this point in question and the one already given is Delta x=0.5\[f(9.5)=f(9+0.5)\approx f(9)+f'(9)(0.5)\]now you have all given information

OpenStudy (anonymous):

do I still need to solve or is that the answer?

OpenStudy (turingtest):

you still need to substitute in the numbers given in the problem and do the simple arithmetic

OpenStudy (turingtest):

what is f(9) ? what is f'(9) ? both are given in your very post

OpenStudy (anonymous):

i got 112.5

OpenStudy (anonymous):

is that right?